OpenStudy (<3pandas18):

Check my answer 3x-3y=15 is the same as y=5-x?

OpenStudy (anonymous):

3x-3y=15 or, 3(x-y)=3*5 or,x-y= 5 or, -y= 5-x or, y= x-5 so the answer is this. r u getting this?

OpenStudy (<3pandas18):

oooooh ok ya i missed that part so my answer would be x-5 instead of 5-x right?

OpenStudy (dinamix):

yup @<3pandas18

OpenStudy (<3pandas18):

thnks :)